For the following system, subtract the second equation from the first equation.
3x - 4y = 7
3x + 2y = -5
What is the resulting equation in simplest form?

Answer by ikleyn(52787)

.
For the following system, subtract the second equation from the first equation.
3x - 4y = 7
3x + 2y = -5
What is the resulting equation in simplest form?
----------------------------------------------------

3x - 4y =  7
3x + 2y = -5

--------------

-4y - 2y = 12,

-6y = 12,

y = -2.